What is the probability of obtaining either a heart or a diamond in a single blind draw from a standard deck of cards?

I really need help to understand how to first set up this problem. I know that there are 13 of each suit. So the hearts would be 13/52 but would the diamonds also be 13/52 or do you subtract 13 from the 52. In other words would the diamonds be 13/39.

Answer by Theo(13342)

either a hear or a diamond would be 13/52 + 13/52 = 26/52
p(a or b) = p(a) + p(b) - p(a and b)
p (a and b) would be the probability of drawing a heart and a diamond with one draw which is impossible, so the probability of p(a and b), in this case, is equal to 0.
you have 13 hearts in the deck and you have 13 diamonds in the deck and you have 0 cards that are both a heart and a diamond.